# The AI Memory Gap: Users Misremember What They Created With AI or Without

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Empirical studies in HCIccs: Computing methodologies Natural language generation
@@ 1. Introduction
This issue matters not only for questions of intellectual ownership, but also because LLMs can shape cognitive processes. For instance, prior research shows that LLMs can influence users’ opinions (Jakesch et al., 2023a), which suggests they might also affect how people remember their own ideas and contributions.
@@
However, we currently lack the fundamental understanding necessary to make informed design choices in this context. In particular, it remains unknown how accurately and confidently people can recall what they created themselves versus what was suggested by AI, and whether this depends on the AI’s role (e.g., idea generation, text composition, or both). This motivates our research questions:
@@
To address these questions, we conducted a within-subjects study with two phases (Figure 1). In Phase 1, participants (N=184) completed ideation and writing tasks either alone (noAI) or with an LLM-based chatbot (withAI). After a one-week delay (Phase 2), they were asked to recall the origin of each idea and elaboration sentence they had produced, specifying whether they had created it themselves or with AI support, and rating their confidence in each decision. To probe memory accuracy more rigorously, we also introduced content not previously seen by participants (“distractors”).
Our results reveal an “AI Memory Gap”, where any AI involvement impaired memory, specifically source memory, and most strongly in mixed human-AI workflows. The odds of correctly attributing the source of ideas were 95% lower when the idea originated from the AI but was elaborated by the human, and 86% lower when the idea was human but the elaboration was generated by the AI, compared to the all-human baseline. When participants consistently used AI in both steps, the interaction partly offset these losses. The same pattern holds true for elaboration source attribution, where using AI for the idea alone reduced the odds of correct attribution by 65%, and using AI for elaboration alone reduced them by 63%. When both idea and elaboration came from AI, the interaction again compensated for the losses, but still remained below the all-human baseline. Confidence mirrored these effects: participants were most confident in the all-human workflow and less confident with AI involvement. Yet, they systematically overestimated their performance, with self-rated accuracy exceeding actual performance by 12% for ideas, and 6% for elaborations.
@@ 2.1.2. Design and perception of writing with AI
The recent survey by Lee et al. (2024) analysed 115 “intelligent” writing tools from papers at HCI and NLP venues, charting a design space with five high-level dimensions. These design choices may impact how people remember their interactions and results. For example, their dimension of “Interaction” includes aspects such as UI layout (separated or integrated AI), interaction metaphor (e.g. tool or agent), and interface paradigm (e.g. chatbot or text editor). For our study, we decided to mimic the currently widespread, de-facto “standard UI” for LLMs: a chatbot in a sidebar.
@@
One notable exception is the HaLLMark system by Hoque et al. (2024), which is explicitly designed for writers to keep a history of prompts and AI contributions throughout their writing process. As with IdeationWeb (Shen et al., 2025) above, this tracks AI use in the UI, whereas we investigate how users remember the source of ideas and texts unaided.
@@ 2.2. Source Memory
Confidence in source decisions is only weakly correlated with accuracy (e.g. see (Horry et al., 2014)). People consistently overestimate their ability to identify the origin of text, especially when they believe that stylistic idiosyncrasies are obvious cues such as in the case of AI writing (Fleckenstein et al., 2024). Metacognitive overconfidence is exacerbated by fluency – the ease with which information is processed – because fluent processing is misinterpreted as evidence of prior self-generation (Huang and Shanks, 2021). LLMs explicitly optimise for fluency, suggesting that AI collaboration may produce particularly compelling illusions of authorship.
@@ 3. Method
To investigate how co-writing with LLMs affects people’s source memory for ideas and elaboration texts, we conducted a two-phase, within-subjects online experiment. Figure 2 gives a visual overview.
Figure 2. Study design and procedure. The experiment consisted of two phases separated by one week. In Phase 1 (top), participants first generated five short ideas (1-3 keywords each) for eight problem statements, alternating between conditions with and without AI assistance (top left). Second, they elaborated on each idea with a one-sentence explanation, again alternating between AI-assisted and unassisted conditions (top right). Problem order and withAI/noAI order were counterbalanced across participants. In Phase 2 (bottom), participants saw 60 elaborations (40 self-generated and 20 distractors) together with the original problem statements. For each, they indicated whether they remembered working on this item, and if so, attributed the source of both the idea and the elaboration (self vs. AI), along with confidence ratings. Distractors included both elaborations from familiar problems (known-topic) and from unseen problems (unknown-topic). This procedure implemented a 2×2 within-subjects design (noAI vs. withAI during ideation and elaboration) to measure item memory and source attribution.
[…]
@@ 3.1. Experiment Design
Our experiment consists of two phases a week apart. In Phase 1, participants created solution ideas and elaboration texts, either with or without AI support. In Phase 2, they indicated the sources of idea and elaboration, given the problem and elaboration text. Participants who completed Phase 1 received an invitation after 6 days to complete Phase 2 within 48 hours. This delay is consistent with source memory studies in the literature (e.g., (Beaufort et al., 2013; Bornstein and Lecompte, 1995; Brown and Halliday, 1991)) and the 48-hour time window was used to reduce drop-out.
@@ 3.1.2. Dependent Variables
In detail, Phase 2 displayed each participant’s own elaboration sentences along with the problem text (40 items). In the same manner, we displayed the distractors (texts not previously seen or generated by the participant, 20 items). For counterbalancing, we randomised the order of problems and solutions for each participant and used this in a “round-robin” procedure (one idea from all problems before returning to another idea from the first problem), thus maximizing the distance between encountering solutions for the same problem. This was motivated by reducing context effects and shortcuts in source attribution.
First, for each elaboration shown in Phase 2, participants indicated whether they remembered working on it (binary: “Yes” / “No”), allowing us to assess item memory performance. Second, they attributed the source of each item’s underlying idea and elaboration (binary: “On my own” / “With AI support”). Third, they rated their confidence in each attribution on a slider ranging from 0 (“Very unsure”) to 100 (“Very sure”).
Finally, at the end of Phase 2, participants estimated their own attribution accuracy for both idea and elaboration source (0–100%) and indicated whether they believed they had more often misattributed content to themselves or to the AI.
@@ 3.2. Participant Recruitment
Our sample comprises 184 generally healthy people. Out of those 89 self-identified as female and 95 as male. All participants were either based in the US (72 people) or the UK (112 people) and all of them listed English as their primary language, with all but 7 participants being native speakers. Their age ranged from 19 to 75 (mean=39.8, SD=12.7). We asked participants how often they use AI tools (e.g., ChatGPT, Claude, Copilot). Most of them (139) reported daily use or several times a week, some (35) once a week or several times a month, and few participants (10) rarely or never user AI. Most participants had a college degree: 82 held a bachelor’s, 28 a master’s, and 7 a doctorate. Eleven reported an associate degree. 37 had a high school diploma and 18 had some college education without a degree. One reported another type of education.
The sample was acquired through the online platform Prolific 111www.prolific.com. We excluded 16 participants from the initially recruited 200 people, as they either did not complete Phase 2 of the experiment within the required three-day window (6 people) or our logs showed that they did not complete the tasks correctly (10 people; e.g., entering “no idea” or always submitting the same text). Participants who completed both parts of the experiment received a total compensation of £23. The average completion time was 107 minutes, resulting in an effective hourly rate of approximately £12.90. The experiment was approved by our institution’s ethics committee.
@@ 3.3.1. Problem Statements
We prepared a pool of ten problem statements for the ideation task in Phase 1, spanning social, technological, and environmental issues. They were designed to spark creative problem-solving without triggering emotional responses or appearing overly opinionated (e.g., “List practical ways a mid-sized company could reduce employee susceptibility to phishing emails.”). Guided by the concept of desirable difficulty (Bjork and Bjork, 2020), we formulated the problems so that generating five ideas would be achievable for everyone within a reasonable amount of time, but challenging enough to encourage effortful cognitive processing. A list of all ten statements can be found in  Section A.1. For each participant, eight of those topics were used in Phase 1, and two served as distractors in Phase 2.
@@ 3.3.2. Distractors
For each problem statement a participant had encountered in Phase 1, we added one additional elaboration sentence from the same problem as a distractor. These known-topic distractors tested whether participants would falsely recognize or misattribute content that was topically familiar but otherwise novel. To reduce the risk that within-topic distractors might be inadvertently similar to participants’ own responses, which could lead to them “correctly” remembering distractors, we chose solution texts that are very distinctive and always included a proper name. For example, for the problem statement “Oceans are heavily polluted with plastic waste. What practical approaches could help reduce plastic pollution in the oceans?”, the within-topic distractor was: “Deploy AquaClean bio-magnetic kelp that’s genetically modified to attract and collect microplastics for easier harvesting.”
@@ 3.3.3. Experimental Software
We developed a custom web application using React  [^2] https://react.dev with Next.js  [^3] https://nextjs.org (Figure 3). It handled communication with a PostgresDB using Prisma [^4] https://www.prisma.io. We used this database to store participant responses as well as some interaction logs. We used the AI-SDK [^5] https://ai-sdk.dev to enable communication between users and the LLM endpoint. We used the OpenAI endpoint “gpt-4.1-mini-2025-04-14” with the system prompt “You are a helpful assistant.” and provided no further context.
For Phase 1, the app showed a welcome screen with study information and consent form, followed by the tasks (Figure 3). In Phase 2, it then displayed the corresponding memory and attribution tasks (Figure 4).
@@
Figure 4. Memory and source attribution in Phase 2: We displayed the task description, the original problem statement, and the solution text created in Phase 1. Participants first had to answer whether they remember working on a solution, if so they had to indicate whether they came up with the underlying idea and the solution text on their own or with AI support. They also had to provide confidence ratings (0-100) for each.
[…]
@@ 3.4. Procedure
All participants completed the study online via Prolific, using a web browser on a desktop or laptop computer. Mobile devices were not permitted. Participants first read an online information sheet and consent form, and if they agreed continued to Phase 1. On average, participants required 74.7 minutes (SD 30.9 minutes) for Phase 1 and 32.8 minutes (SD 16.6 minutes) for Phase 2. After completing Phase 1, they started Phase 2 after 6 days and 5 hours (SD 8 hours 46 minutes) on average.

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3791494isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Empirical studies in HCIccs: Computing methodologies Natural language generation
@@ 1. Introduction
This issue matters not only for questions of intellectual ownership, but also because AI can shape cognitive processes. For instance, prior research shows that AI-edited images and videos can induce false recollections (Pataranutaporn et al., 2025). Similarly, co-creating or revising text with AI may blur the line between own experience and AI-generated modifications, potentially influencing source memory. Furthermore, Jakesch et al. (2023a) found that LLMs can influence users’ opinions, which suggests they might also affect how people remember their own ideas and contributions.
@@
As AI tools and their capabilities evolve, the need to track and disclose AI involvement will remain a recurring challenge. This prompts institutions to frequently update their policies, requiring post hoc declarations of when and where AI was used. For example, academic journals may mandate disclosure of AI assistance for manuscripts drafted before such rules were in place (e.g., Elsevier’s update on GenAI use [^1] https://www.elsevier.com/about/policies-and-standards/generative-ai-policies-for-journals), and in the legal domain, courts may require practitioners to specify which parts of a document were generated by AI, sometimes months after the original filing (cf. “Mata v. Avianca, Inc.” (Lyon, 2023)).
Despite these issues, we currently lack the fundamental understanding necessary to make informed design choices in this context. In particular, it remains unknown how accurately and confidently people can recall what they created themselves versus what was suggested by AI, and whether this depends on the AI’s role (e.g., idea generation, text composition, or both). Thus, our study focuses on ideation and elaboration as fundamental elements of writing workflows. Separating these stages also allows us to examine how using AI at different points in the writing process affects people’s memory. This motivates our research questions:
@@
To address these questions, we conducted a within-subjects study with two phases (Fig. 1). In Phase 1, participants (N=184) completed ideation and writing tasks either alone (noAI) or with an LLM-based chatbot (withAI). After a one-week delay (Phase 2), they were asked to recall the origin of each idea and elaboration sentence they had produced, specifying whether they had created it themselves or with AI support, and rating their confidence in each decision. To probe memory accuracy more rigorously, we also introduced content not previously seen by participants (“distractors”).
Our results reveal an “AI Memory Gap”, where any AI involvement impaired memory, specifically source memory, and most strongly in mixed human-AI workflows. The odds of correctly attributing the source of ideas were 95% lower when the idea originated from AI but was elaborated by the human, and 86% lower when the idea was human but the elaboration was generated by AI, compared to the all-human baseline. When participants consistently used AI in both steps, the interaction partly offset these losses. The same pattern holds true for elaboration source attribution, where using AI for the idea alone reduced the odds of correct attribution by 65%, and using AI for elaboration alone reduced them by 63%. When both idea and elaboration came from AI, the interaction again partly compensated for the losses, but still remained below the all-human baseline. Confidence mirrored these effects: Participants were most confident in the all-human workflow and less confident with AI involvement. Yet, they systematically overestimated their performance, with self-rated accuracy exceeding actual performance by 12% for ideas, and 6% for elaborations.
@@ 2.1.2. Design and perception of writing with AI
The recent survey by Lee et al. (2024) analysed 115 “intelligent” writing tools from papers at HCI and NLP venues, charting a design space with five high-level dimensions. These design choices may impact how people remember their interactions and results. For example, their dimension of “Interaction” includes aspects such as UI layout (separated or integrated AI), interaction metaphor (e.g. tool or agent), and interface paradigm (e.g. chatbot or text editor). For our study, we decided to replicate a common AI-assisted writing workflow: using an AI chatbot alongside a text editor. To keep both visible without switching apps/tabs, we embedded the chatbot into a sidebar within our UI, similar to placing a text editor (e.g., Word, Docs) and AI tool (e.g., ChatGPT, Claude) side by side.
@@
One notable exception is the HaLLMark system by Hoque et al. (2024), which is explicitly designed for writers to keep a history of prompts and AI contributions throughout their writing process. As with IdeationWeb (Shen et al., 2025) above, this tracks AI use in the UI.
Finally, earlier work on computer-supported collaborative work (CSCW) tracked and visualised human contributions (e.g. (Bergstrom and Karahalios, 2007; Shi et al., 2023)). While some solutions for detecting and tracking AI use exist in research and products (e.g. (S. Gehrmann, H. Strobelt, and A. Rush (2019); M. N. Hoque, T. Mashiat, B. Ghai, C. D. Shelton, F. Chevalier, K. Kraus, and N. Elmqvist (2024); 31)), our findings motivate revisiting the CSCW design space more broadly – now for tracking AI. Some adaptations might be necessary: In contrast to human-human collaboration, there is no other person that might keep own notes or remember their contribution independent of what is tracked by the system, and people tend to give less credit to AI than a human writer (Draxler et al., 2024). Overall, we see our contribution as complementary to system/design work, in that we quantify an unaided human memory baseline. For example, HaLLMark (Hoque et al., 2024) could not relate to memory at all in its design rationale, and framed it around accountability, ownership, and external policies. Our findings provide an important “internal” addition to this set: imperfect human memory of interaction with AI.
@@ 2.2. Source Memory
After relating our work to the literature on writing with AI, we now introduce the concept of source memory to contextualise these insights within established concepts from memory research.
@@
Confidence in source decisions is only weakly correlated with accuracy (e.g. see (Horry et al., 2014)). People consistently overestimate their ability to identify the origin of text, especially when they believe that stylistic idiosyncrasies are obvious cues such as in the case of AI writing (Fleckenstein et al., 2024). Metacognitive overconfidence is exacerbated by fluency – the ease with which information is processed – as fluent processing is misinterpreted as evidence of prior self-generation (Huang and Shanks, 2021). LLMs explicitly optimise for fluency, suggesting that AI collaboration may produce particularly compelling illusions of authorship.
@@ 3. Method
To investigate how co-writing with LLMs affects people’s source memory for ideas and elaboration texts, we conducted a two-phase, within-subjects online experiment. Fig. 2 gives a visual overview.
Figure 2. Study design and procedure. The experiment consisted of two phases separated by one week. In Phase 1 (top), participants first generated five short ideas (1-3 keywords each) for eight problem statements, alternating between conditions with and without AI assistance (top left). Second, they elaborated on each idea with a one-sentence explanation, again alternating between AI-assisted and unassisted conditions (top right). Problem order and withAI/noAI order were counterbalanced across participants. Phase 2 (bottom) started after around one week (accessible after 6 days with a 48-hour completion window). In Phase 2, participants saw 60 elaborations (40 self-generated and 20 distractors) together with the original problem statements. For each, they indicated whether they remembered working on this item, and if so, attributed the source of both the idea and the elaboration (self vs. AI), along with confidence ratings. Distractors included both elaborations from familiar problems (known-topic) and from unseen problems (unknown-topic). This procedure implemented a 2×2 within-subjects design (noAI vs. withAI during ideation and elaboration) to measure item memory and source attribution.
@@ 3.1. Experiment Design
Our experiment consists of two phases a week apart. In Phase 1, participants created solution ideas and elaboration texts, either with or without AI support. We decided to separate the writing task into two fundamental steps, ideation and elaboration, instead of letting participants solve the prescribed problems in one go. This structured approach made it possible to more clearly isolate the effects of using AI during idea generation versus during sentence-level writing. In Phase 2, they indicated the sources of idea and elaboration, given the problem and elaboration text.
Participants who completed Phase 1 received an invitation after 6 days to complete Phase 2 within 48 hours. This delay is consistent with source memory studies in the literature (e.g., (Beaufort et al., July 2013; Bornstein and Lecompte, 1995; Brown and Halliday, 1991)) and the 48-hour time window was used to reduce drop-out.
@@ 3.1.1. Independent Variables
[…]
Table 1 shows example inputs from the participants to illustrate the experimental process.
Topic/Problem Idea Idea source Elaboration Elaboration source
Think of features that could help blind or low-vision people use a city’s bus/train app more easily. Live agent, travel guidance noAI The app could offer an option to connect with a live agent who can provide real-time travel assistance, directions, and updates to support blind or low-vision users during their commute. noAI
Climate change is causing more frequent and severe weather events. What innovative ideas could help communities better prepare for and adapt to these changes? Resilient Infrastructure withAI When there’s resilient infrastructures, there are top and tough buildings, roads, and utilities that can withstand extreme weather like floods, storms, and heatwaves. They reduce damage costs, keep essential services running and faster recoveries should any disaster happen noAI
Come up with ways students can fairly distribute work for an essay in a group project. each writes a paragraph noAI Each student takes responsibility for writing one paragraph, ensuring equal contribution and allowing the group to combine their individual efforts into a cohesive essay. withAI
Oceans are heavily polluted with plastic waste. What practical approaches could help reduce plastic pollution in the oceans? Promote eco-friendly materials withAI Promoting the use of eco-friendly materials can significantly reduce plastic waste and help protect our oceans from pollution. withAI
Table 1. Example ideas and corresponding elaborations created by participants during the study.
This table lists four example topics used in the study, along with participants’ inputs, and the sources of those ideas and elaborations.
@@ 3.1.2. Dependent Variables
In detail, Phase 2 presented 40 items, where each item consisted of the original problem text paired with the participant’s own elaboration sentence, which they created in Phase 1. The underlying ideas were not shown again. In addition, we presented 20 distractor items, each of which consisted of a problem text paired with an elaboration the participant had not previously generated or seen. For counterbalancing, we randomised the order of problems and solutions for each participant and used this in a “round-robin” procedure (one idea from all problems before returning to another idea from the first problem), thus maximizing the distance between encountering solutions for the same problem. This was motivated by reducing context effects and shortcuts in source attribution.
Concretely, for each elaboration shown in Phase 2, we first asked participants, “Do you remember working on this solution? (binary: “Yes” / “No”), allowing us to assess item memory performance. Second , we asked “Did you come up with the underlying idea for the solution on your own or with AI-support?” and “Did you write this solution text on your own or with AI-support? (binary: “On my own” / “With AI support”), to measure source memory performance. In addition, for each attribution, they rated their confidence on a slider ranging from 0 (“Very unsure”) to 100 (“Very sure”). Finally, at the end of Phase 2, participants estimated their own attribution accuracy for both idea and elaboration source. They could answer using a slider ranging from 0% (“None”) to 100% (“All”), and indicated whether they believed they had more often misattributed content to themselves or to the AI.
@@ 3.2. Participant Recruitment
Our sample consisted of 184 healthy adults. Out of those, 89 self-identified as female and 95 as male. All participants were either based in the US (72 people) or the UK (112 people) and all of them listed English as their primary language, with all but 7 participants being native speakers. Their age ranged from 19 to 75 (mean=39.8, SD=12.7). We asked participants how often they use AI tools (e.g., ChatGPT, Claude, Copilot). Most of them (139) reported daily use or several times a week, some (35) once a week or several times a month, and few participants (10) rarely or never use AI. Most participants had a college degree: 82 held a bachelor’s, 28 a master’s, and 7 a doctorate. Eleven reported an associate degree. 37 had a high school diploma and 18 had some college education without a degree. One reported another type of education.
[…]
@@ 3.3.1. Problem Statements
We prepared a pool of ten problem statements for the ideation task in Phase 1, spanning social, technological, and environmental issues. We designed them to spark creative problem-solving without triggering emotional responses or appearing overly opinionated (e.g., “List practical ways a mid-sized company could reduce employee susceptibility to phishing emails.”). Guided by the concept of desirable difficulty (Bjork and Bjork, 2020), we formulated the problems so that generating five ideas would be achievable for everyone within a reasonable amount of time, but challenging enough to encourage effortful cognitive processing. A list of all ten statements can be found in Section A.1. For each participant, eight of those topics were used in Phase 1, and two served as distractors in Phase 2.
@@ 3.3.2. Distractors
For each problem statement a participant had encountered in Phase 1, we added one additional elaboration sentence from the same problem as a distractor. These known-topic distractors tested whether participants would falsely recognize or misattribute content that was topically familiar but otherwise novel. To reduce the risk that these within-topic distractors might be inadvertently similar to participants’ own responses, which could lead to them “correctly” remembering distractors, we chose solution texts that are distinctive and always included a proper name, while still being plausible within the topic. Although this choice may have made these distractors comparatively easy to reject, their main purpose was to prevent participants from making decisions based solely on general topic or problem familiarity. They also provided a way to assess false recognition.
For example, for the problem statement “Oceans are heavily polluted with plastic waste. What practical approaches could help reduce plastic pollution in the oceans?”, the within-topic distractor was: “Deploy AquaClean bio-magnetic kelp that’s genetically modified to attract and collect microplastics for easier harvesting.”
@@ 3.3.3. Experimental Software
We developed a custom web application using React [^3] https://react.dev with Next.js [^4] https://nextjs.org (Fig. 3). It handled communication with a PostgresDB using Prisma [^5] https://www.prisma.io. We used this database to store participant responses as well as interaction logs. While we did not record any interaction outside of our experimental software (e.g., with external AI tools), we logged “focus lost” events (e.g., when particpants switched to another tab during the study) and selected keystrokes (e.g., “paste attempts” in the noAI condition) to verify integrity. We used the AI-SDK [^6] https://ai-sdk.dev to enable communication between users and the LLM endpoint. We used the OpenAI endpoint “gpt-4.1-mini-2025-04-14” with the system prompt “You are a helpful assistant.” and provided no further context.
For Phase 1, the app showed a welcome screen with study information and consent form, followed by the tasks (Fig. 3). In Phase 2, it then displayed the corresponding memory and attribution tasks (Fig. 4).
@@
Figure 4. Memory and source attribution in Phase 2: We displayed task description, original problem statement, and solution text created in Phase 1. Participants first answered whether they remember working on a solution. If so, they indicated whether they came up with the underlying idea and the solution text on their own or with AI support. They also provided confidence ratings (0-100) for each.
@@ 3.4. Procedure
All participants completed the study online via Prolific, using a web browser on a desktop or laptop computer. Mobile devices were not permitted. In addition, our task explicitly asked participants to use the AI chat within our experimental software (withAI) or to solve the task on their own (noAI). We found no differences between the conditions in the number of “focus lost” events over time or other suspicious activity. This indicates that participants generally adhered to these instructions.
Participants first read an online information sheet and consent form, and if they agreed continued to Phase 1. On average, participants required 74.7 minutes (SD 30.9 minutes) for Phase 1 and 32.8 minutes (SD 16.6 minutes) for Phase 2. After completing Phase 1, they started Phase 2 after 6 days and 5 hours (SD 8 hours 46 minutes) on average.